OpenStudy (anonymous):

Help!! Ill give a medal to whoever answers!! If f(x) = x2 + 3, find f(5). 13 22 28 4

OpenStudy (nurali):

f(x) = x2 + 3 f(5)=(5)2+3 f(5)=25+3 f(5)=28

OpenStudy (anonymous):

Thank you! @Nurali